# biophysics, medical physics, biomechanics

### v1

@@ Abstract
Airway constriction and blockage in obstructive lung diseases create barriers to effective drug deposition by altering how different regions of the lungs are ventilated. Established computational particle deposition models have not accounted for these impacts of disease. We present a new particle deposition model that calculates ventilation based on the resistance of each airway, such that ventilation patterns respond to airway constriction. We incorporate distal airway constrictions representative of cystic fibrosis, and assess the resulting impact on deposition down to the single-airway scale. We demonstrate how constriction reduces deposition in the airways directly distal and proximal to the affected airways. When multiple constrictions are clustered together, deposition in the central airways proximal to the constrictions is more strongly reduced, and deposition in the other central airways is generally increased. This results in more uneven deposition in both the central and distal airways, even when constrictions affect only the distal airways. We use our model to calculate lung clearance index (LCI), a clinical measure of ventilation heterogeneity, in lungs with constrictions of various severities localised to one lobe. We find an increase in LCI coinciding with a significant drop in deposition throughout the affected lobe.
@@ 1 Introduction
Inhaled therapies, including mucolytics and antibiotics, are commonly used to treat cystic fibrosis (CF) [henke2007mucolytics, smith2022inhaled]. Recent evidence has also shown that inhaled gene therapies can also halt or slow decline of lung function [alton2015repeated]. Airway constriction or blockage caused by thickened mucus layers or mucus plugging can create barriers to the effective application of inhaled therapies in CF. Typically, it is desired for inhaled drugs to be deposited evenly throughout the lungs’ conducting airways, including in the distal conducting airways. These small airways can be susceptible to constriction or blockage even in early disease [tiddens2010cystic], so delivering a sufficient dose to them, particularly in more diseased and poorly ventilated regions of the lungs, can be challenging. Barriers to effective application of inhaled therapies are not unique to CF; airway blockage can also alter patterns of inhaled particle deposition in chronic obstructive pulmonary disease (COPD) or severe asthma [darquenne_2012_aerosol, labiris2003pulmonary]. Experimental studies have shown increased particle deposition in obstructive lung diseases, potentially caused by increased flow rates through narrowed airways [anderson1989effect], but measuring how airway constriction affects the detailed spatial patterns of deposition is difficult. Computational particle-deposition modelling has the potential to provide detailed spatial information but, to do so, the heterogeneous ventilation induced by airway constriction must be accounted for.
@@
Other studies have used just CFD to model particle deposition (see, e.g., the review by Longest & Holbrook [longest2012silico]). This is typically limited to modelling a relatively small number of airways. Some attempts have been made to incorporate bronchoconstriction in these models, but they have largely focused on simulating particle deposition in only two or three generations of airways [longest2006transport, farkas2007simulation]. Walenga & Longest [walenga2016current] explored the effects of airway constriction in a geometry representative of the whole lungs, in which each lobe was modelled by a single path of airways, with the other airways branching off that single path being truncated. They applied a uniform constriction to all airways in their model; thorough exploration of heterogeneous patterns of distal airway constriction are beyond the current capabilities of CFD models.
@@
In this study, we present a whole-lung particle-deposition model, in which ventilation of the lungs is derived based on the resistance of each conducting airway and is driven by the expansion and contraction of individual acini. Transport of an inhaled gas and deposition of particles from that gas throughout the lungs are then calculated. An overview of the model setup is provided in §2, with full details in the supplementary material. The model can predict significant changes to patterns of ventilation and particle deposition throughout the lungs when there is bronchoconstriction. We investigate these effects by applying spatially heterogeneous patterns of distal airway constrictions, representative of airway disease. We apply several qualitatively different patterns of constriction, including constricting airways in one or two lobes only, constricting a number of localised clusters of distal airways, or constricting a number of airways distributed randomly throughout the lungs. We simulate associated clinical measures of ventilation heterogeneity (lung clearance index, LCI) and regional dosing (gamma scintigraphy).
@@ 2.1 Lung networks
To enable direct comparison of our model with MPPD, and to minimise differences purely due to lung size, we rescale the four lung geometries so that they are representative of an adult’s lungs with functional residual capacity (FRC) of $3.3\mathrm{L}$. Given the measured FRC, $V_{\rm FRC}$, from each geometry, we uniformly increase the length and radius of each airway in that geometry by a factor of $(3.3\mathrm{L}/V_{\rm FRC})^{1/3}$. This results in a range of airway dead-space volumes between $101\mathrm{mL}$ and $145\mathrm{mL}$, which is realistic for adult lungs, noting we do not include the oral cavity [hart1963relation]. In figure 2, we present data from simulations in all four geometries and compare these to MPPD to validate the model. Subsequently, we present results from simulations in one of the four geometries, chosen arbitrarily, to focus on comparing deposition in the unconstricted geometry with deposition after applying constrictions to the distal airways.
@@ 2.2 Ventilation
We assume that particle transport and deposition do not affect ventilation, so we first simulate ventilation and then use the calculated flow rates in each airway to simulate particle transport and deposition. The ventilation model is based on a previous model from Whitfield et al [whitfield2020spectral]. We provide an overview here and details in the supplementary material, §S2. We represent each lung geometry as a network of edges and vertices embedded in 3D space, with each edge representing an airway and vertices being placed at the ends of each edge. The resistance to air flow of each edge is calculated via Poiseuille’s law, providing a linear relation between the difference in pressure along each airway and the flow rate through it. Whilst Poiseuille’s law is not likely to accurately capture the complex dynamics in the large airways, it provides a computationally efficient simple approximation. We have tested a nonlinear resistance model that accounts for the formation of flow boundary layers at bifurcations [pedley1970prediction], but found it had minimal impact on total deposition.
@@ 2.3 Particle transport and deposition
We define $s$ in (1) by assuming that deposition occurs via three mechanisms: inertial impaction, gravitational sedimentation and diffusion. Following the approach of many other studies [hofmann2011modelling], we assume these mechanisms act independently, and approximate deposition rates using semi-empirical or derived formulae. For impaction, we use a formula from Zhang et al. [zhang_1997_impaction] derived from CFD simulations in airway bifurcation geometries. For sedimentation, we use a formula from Pich [pich_1972_theory], which was also used by Oakes et al. [oakes_aerosol_2017]. For diffusion, we use a formula from Ingham [ingham1991diffusion], which has been validated against experimental results and CFD for sub-micron nanoparticles [longest2007computational]. We do not model extrathoracic deposition, which is typically significant primarily for particles at least $6\mu\mathrm{m}$ in diameter [darquenne_2016_total], focusing instead on patterns of lung deposition. We mostly present simulations of $4\mu\mathrm{m}$ diameter particles, as this is representative of typical median particle sizes generated by nebulisers used to administer inhaled therapeutics [collins2009nebulizer].
@@
We model particle deposition during a single breath, assuming that the initial concentration of particles is zero everywhere, and that the flux of concentration into the trachea is fixed during inhalation. To solve the discrete advection-diffusion equation, we approximate the time derivative using a first-order backwards Euler finite-difference scheme, and solve the resulting system of equations using the BiCGSTAB solver in the Eigen C++ library [eigen_2010_v3]. We use a time step of $\Delta t=0.01\mathrm{s}$, which is small enough that total deposition is independent of the exact value. To discretise the airways to solve (1), we ensure all edges within each airway have the same length, every airway, including in the acini, contains at least eight edges and edges are at most $200\mu\mathrm{m}$ long. The four discretised lung networks have between 2.5 and 3 million edges each. This discretisation ensures convergence of both conducting and acinar airway deposition; in the simulation presented in figure 2(a), approximately doubling the total number of edges changes the total deposition by only $0.2\%$. One typical simulation takes around two hours on a single-core $3.0\mathrm{GHz}$ processor.
@@ 2.4 Application of airway constrictions
2.5 Multiple-breath washout and gamma scintigraphy
@@
We use particle-deposition predictions to simulate gamma scintigraphy, a clinical measurement of regional dosing. To do so, we divide the 2D plane into many pixels, sum the deposition that occurs within each pixel (across the whole depth of the lungs), and use kernel density estimation in Matlab to generate a deposition distribution that simulates a scintigraphy image. Further details of this process are in the supplementary material, §S5.
@@ 3.1 Particle deposition in the lungs without airway constriction
Figure 2: Data from simulations in the unconstricted lung geometries. In (a)-(d), particle diameter is $4\mu\mathrm{m}$. (a) Deposited fraction within each conducting airway in one geometry, shown as a planar graph. Deposition at bifurcations is assigned to the parent airway. (b) Simulated scintigraphy plot for the same example. (c) Deposited fraction in each lobe, showing mean $\pm$ one standard deviation across the four geometries. (d) Deposited fraction in each generation, showing mean $\pm$ one standard deviation across the four geometries and comparison with MPPD. (e) Total and conducting-airways deposition, versus particle diameter.
@@
To validate the model in healthy lung geometries, we compare deposition in the four lung geometries against results from MPPD (figure 2d,e). Deposition in each airway generation shows generally good agreement for $4\mu\mathrm{m}$ diameter particles (figure 2d), despite MPPD using a lung geometry based on different morphometric data. There is also good agreement with MPPD predictions of total deposition over a range of particle sizes, with only minor quantitative differences for particle diameters smaller than $0.1\mu\mathrm{m}$ or larger than $1\mu{\mathrm{m}}$. Higher conducting-airways deposition for large particles in our model is likely due to the central airways being narrower on average than in the MPPD lung geometry, causing higher impaction. We use different semi-empirical formulae than in MPPD to determine deposition rates, but despite this, agreement is still good. The MPPD model has been validated against experimental data [asgharian_2022_mixing], so good agreement here suggests our model accurately captures deposition fractions in healthy lungs.
@@ 3.2 Effects of varying airway constriction severity
Figure 3: Simulations in which all generation 12-15 airways in the left upper lobe are constricted with varying severities (fraction by which airway radii are reduced). In each simulation, the constriction severity is the same for all left upper lobe airways. (a) Change in deposition in the affected lobe (LU), and the other lobes, separated into central conducting-airways deposition, distal conducting-airways deposition and acinar deposition. (b) LCI, from MBW simulations in the same constricted geometries. (c) Change in individual airway deposition in the simulation with constriction severity 0.825.
@@
To explore how the severity of airway constriction impacts deposition, we first constrict all generation 12-15 airways in a single lobe, with a range of constriction severities. The effect of increasing constriction severity on deposition in the affected airways is non-monotonic (figure 3a). Distal-airways deposition is not significantly impacted by mild constrictions but, when the severity exceeds 0.4, it increases as impaction is significantly enhanced in the narrowed airways. As the severity is increased further beyond 0.675, the increased resistance of the affected airways causes flow to be reduced into the affected lobe to such a degree that distal deposition decreases rapidly. Increased impaction in narrowed airways has been highlighted previously in simpler models [martonen2003silico], but these have not captured the subsequent decrease as airway resistance is further increased and ventilation patterns are altered. Central and acinar deposition in the affected lobe decrease as the constriction severity is increased and ventilation of the lobe is reduced (figure 3a). In the other lobes, deposition increases due to a higher proportion of the inhaled particles entering these lobes and the speed of air flow through them being faster. The assumption of fixed tidal volume means that these local changes in flow rate are solely due to redistribution of lobar ventilation. The strongest increase in deposition in the unconstricted lobes is in the central airways since impaction is the dominant deposition mechanism there, and this is enhanced by faster flow. By contrast, in the distal conducting airways, where sedimentation is the dominant mechanism, there is minimal change to deposition as shorter residence times for particles leads to reduced sedimentation, enough to largely offset any increased impaction.

### v3

@@ Abstract
Airway constriction and blockage in obstructive lung diseases cause ventilation heterogeneity and create barriers to effective drug deposition. Established computational particle-deposition models have not accounted for these impacts of disease. We present a new particle-deposition model that calculates ventilation based on the resistance of each airway, such that ventilation responds to airway constriction. The model incorporates distal airway constrictions representative of cystic fibrosis, allowing us to investigate the resulting impact on patterns of deposition. Unlike previous models, our model predicts how constrictions affect deposition in airways throughout the lungs, not just in the constricted airways. Deposition is reduced in airways directly distal and proximal to constrictions. When constrictions are clustered together, central-airways deposition can increase significantly in regions away from constrictions, but distal-airways deposition in those regions remains largely unchanged. We use our model to calculate lung clearance index (LCI), a clinical measure of ventilation heterogeneity, after applying constrictions of varying severities in one lobe. We find an increase in LCI coinciding with significantly reduced deposition in the affected lobe. Our results show how the model provides a framework for development of computational tools that capture the impacts of airway disease, which could significantly affect predictions of regional dosing.
@@ 1 Introduction
Inhaled therapies, including mucolytics and antibiotics, are commonly used to treat cystic fibrosis (CF) [henke2007mucolytics, smith2022inhaled]. Recent evidence has also shown that inhaled gene therapies can halt or slow decline of lung function [alton2015repeated]. Airway constriction or blockage caused by thickened mucus layers or mucus plugging can create barriers to the effective application of inhaled therapies in CF. Typically, it is desired for inhaled drugs to be deposited evenly throughout the lungs’ conducting airways, including in the distal conducting airways. These small airways can be susceptible to constriction or blockage even in early disease [tiddens2010cystic], so delivering a sufficient dose to them, particularly in more diseased and poorly ventilated regions of the lungs, can be challenging. Barriers to effective application of inhaled therapies are not unique to CF; airway blockage can also alter patterns of inhaled particle deposition in chronic obstructive pulmonary disease (COPD) or severe asthma [darquenne_2012_aerosol, labiris2003pulmonary]. Experimental studies have shown that, whilst deposition is evidently reduced in blocked airways due to a lack of ventilation through them, total lung deposition can be increased in patients with obstructive lung disease [anderson1989effect]. Innovation in the design of inhalers and nebulisers, and how they are administered, can make drug delivery to the lungs more efficient [laube2000targeting, de2015effects], but ventilation heterogeneity in patients
with obstructive lung disease is still a significant barrier to uniform drug deposition throughout the lungs, particularly in the diseased airways [venegas2005self, venegas2023measuring]. Experimental techniques, such as gamma scintigraphy, can provide information on regional particle deposition in patients, including measures such as central-to-peripheral deposition ratio [de2010lung, virchow2018lung]. Resolution is limited in imaging, and whilst new technologies are providing more detailed data on regional lung deposition, resolving particle deposition on the scale of individual small airways is beyond the capabilities of current experimental measurements. Computational particle-deposition modelling has the potential to provide detailed spatial information on patterns of local particle deposition in the small airways but, to do so, the heterogeneous ventilation induced by airway constriction must be accounted for.
@@
Other studies have used CFD to model particle deposition (see, e.g., the review by Longest & Holbrook [longest2012silico]). This is typically limited to modelling a relatively small number of airways, and whilst relative ventilation of the lungs’ five lobes may be inferred from patient-specific imaging in some models [de2010validation, hajian2016functional], the small airways are not generally modelled, exhalation is generally not captured, and neither is any ventilation heterogeneity beyond differences in lobar ventilation. Some attempts have been made to incorporate bronchoconstriction in CFD models, but they have largely focused on simulating particle deposition in only two or three generations of airways [longest2006transport, farkas2007simulation]. Walenga & Longest [walenga2016current] explored the effects of airway constriction in a geometry representative of the whole lungs, in which each lobe was modelled by a single path of airways, with the other airways branching off that single path being truncated. They applied a uniform constriction to all airways in their model; thorough exploration of heterogeneous patterns of distal airway constriction are beyond the current capabilities of CFD models.
@@
In this study, we present a whole-lung particle-deposition model, in which ventilation of the lungs is derived based on the resistance of each conducting airway and is driven by the expansion and contraction of individual acini. Transport of an inhaled gas and deposition of particles from that gas throughout the lungs are then calculated. Unlike models such as MPPD [asgharian2001particle] that assume uniform ventilation, our model is capable of making physics-based predictions of how airway constrictions throughout the lungs lead to non-uniform, heterogeneous ventilation and the impacts this has on patterns of particle deposition. This advancement in modelling inhaled drug delivery in diseased lungs lays the groundwork for future development of patient-specific particle deposition models that can capture the effects of airway disease. In this study, we demonstrate how the predictions of particle deposition by the model can be fundamentally altered by airway constriction; the predictions provide understanding and quantification of the physical mechanisms by which airway constriction and blockage impact patterns of deposition throughout the lungs. By simulating particle deposition in lungs in which patterns of small airway constriction representative of airway diseases such as CF have been applied, we establish that our model can make predictions that have biologic plausiblity, which match with existing understanding of how airway blockage reduces deposition to those blocked airways, and which provide detailed insight into how deposition in each individual airway in the lungs may be affected by localised airway constrictions.
An overview of the model setup is provided in §2, with full details in the supplementary material. The model can predict significant changes to patterns of ventilation and particle deposition throughout the lungs when there is bronchoconstriction. We investigate these effects by applying spatially heterogeneous patterns of distal airway constrictions, representative of airway disease. We apply several qualitatively different patterns of constriction, including constricting airways in one or two lobes only, constricting a number of localised clusters of distal airways, or constricting a number of airways distributed randomly throughout the lungs. We simulate associated clinical measures of ventilation heterogeneity (lung clearance index, LCI) and regional dosing (gamma scintigraphy).
@@ 2.1 Lung networks
To enable direct comparison of our model with MPPD, and to minimise differences purely due to lung size, we rescale the four lung geometries so that they are representative of an adult’s lungs with functional residual capacity (FRC) of $3.3\mathrm{L}$. Given the measured FRC, $V_{\rm FRC}$, from each geometry, we uniformly increase the length and radius of each airway in that geometry by a factor of $(3.3\mathrm{L}/V_{\rm FRC})^{1/3}$. This results in a range of airway dead-space volumes between $101\mathrm{mL}$ and $145\mathrm{mL}$, which is realistic for adult lungs, noting we do not include the oral cavity [hart1963relation]. The original FRC values were between $1.1\mathrm{L}$ and $1.9\mathrm{L}$. We focus attention on simulating deposition in the scaled lung geometries so that we can compare directly to MPPD. We do not present results from simulations in the smaller, unscaled geometries, but we found that the main qualitative difference was higher inertial impaction in the central airways in the unscaled geometries, compared to in the scaled geometries, since the airway radii are smaller. Increased deposition in the central airways of children compared to adults has been highlighted in previous modelling studies (e.g., [poorbahrami2021whole]). Investigating patterns of deposition in children or adolescents with CF is beyond the scope of this study. In figure 2, we present data from simulations in all four scaled geometries and compare these to MPPD to validate the model. Subsequently, we present results from simulations in one of the four geometries, chosen arbitrarily, to focus on comparing deposition in the unconstricted geometry with deposition after applying constrictions to the distal airways.
@@ 2.2 Ventilation
We assume that particle transport and deposition do not affect ventilation, so we first simulate ventilation and then use the calculated flow rates in each airway to simulate particle transport and deposition. The ventilation model is based on a previous model from Whitfield et al. [whitfield2020spectral]. We provide an overview here and details in the supplementary material, §S2. We represent each lung geometry as a network of edges and vertices embedded in 3D space, with each edge representing an airway and vertices being placed at the ends of each edge. The resistance to air flow of each edge is calculated via Poiseuille’s law, providing a linear relation between the difference in pressure along each airway and the flow rate through it. Whilst Poiseuille’s law is not likely to accurately capture the complex dynamics in the largest central airways, it provides a computationally efficient simple approximation. The underlying assumption in using Poiseuille’s law is that flow in the airways is laminar, which is indeed the case in the smaller airways. However, in the largest central airways, there is likely to be turbulent flow, which is not captured by Poiseuille’s law. We have also tested the nonlinear resistance model due to Pedley et al. [pedley1970prediction], which accounts for the formation of flow boundary layers at bifurcations, in simulations in healthy lung geometries but found it had minimal impact on total deposition.
@@ 2.3 Particle transport and deposition
We define $s$ in (1) by assuming that deposition occurs via three mechanisms: inertial impaction, gravitational sedimentation and diffusion. We mostly present simulations of $4\mu\mathrm{m}$ diameter particles, as this is representative of the typical mass median aerodynamic diameter (MMAD) of particles generated by nebulisers used to administer inhaled therapeutics [collins2009nebulizer]. For $4\mu\mathrm{m}$ diameter particles, the dominant deposition mechanisms are likely to be impaction and sedimentation [hofmann2011modelling]. We also include deposition by diffusion in the model, which enables the model to describe deposition of smaller particles also. In figure 2 below, we validate predictions against MPPD for a range of particle diameters from approximately $10^{-2}\mu\mathrm{m}$ to $8\mu\mathrm{m}$. Following the approach of many other studies [hofmann2011modelling], we assume that the three deposition mechanisms act independently, and approximate deposition rates using semi-empirical or derived formulae. For impaction, we use a formula from Zhang et al. [zhang_1997_impaction] derived from CFD simulations in airway bifurcation geometries. For sedimentation, we use a formula from Pich [pich_1972_theory], which was also used by Oakes et al. [oakes_aerosol_2017]. For diffusion, we use a formula from Ingham [ingham1991diffusion], which has been validated against experimental results and CFD for sub-micron nanoparticles [longest2007computational]. We do not model extrathoracic deposition, which is typically significant primarily for particles at least $6\mu\mathrm{m}$ in diameter [darquenne_2016_total], focusing instead on patterns of lung deposition.
@@
We model particle deposition during a single breath, assuming that the initial concentration of particles is zero everywhere, and that the flux of concentration into the trachea is fixed during inhalation. To solve the discrete advection-diffusion equation, we approximate the time derivative using a first-order backwards Euler finite-difference scheme, and solve the resulting system of equations using the BiCGSTAB solver in the Eigen C++ library [eigen_2010_v3]. Using an implicit, backwards finite difference scheme provides improved stability of numerical solutions compared to many explicit schemes. We use a time step of $\Delta t=0.01\mathrm{s}$, which is small enough that total deposition is independent of the exact value. To discretise the airways to solve (1), we ensure all edges within each airway have the same length, every airway, including in the acini, contains at least eight edges and edges are at most $200\mu\mathrm{m}$ long. The four discretised lung networks have between 2.5 and 3 million edges each. This discretisation ensures convergence of both conducting and acinar airway deposition; in the simulation presented in figure 2(a), approximately doubling the total number of edges changes the total deposition by only $0.2\%$. One typical simulation takes around two hours on a single-core $3.0\mathrm{GHz}$ processor.
@@ 2.4 Application of airway constrictions
2.5 Simulated multiple-breath washout and gamma scintigraphy
@@
We use particle-deposition predictions to simulate gamma scintigraphy, a clinical measurement of regional dosing, which uses radiolabelling to image the distribution of particles in real patients. To do so, we divide the 2D plane into many pixels, sum the deposition that occurs within each pixel (across the whole depth of the lungs), and use kernel density estimation in Matlab to generate a deposition distribution that mimics a scintigraphy image. The simulated scintigraphy is therefore derived entirely from the particle-deposition calculations, and is simply a method for visualising the simulation outcomes. Further details of the process of generating a simluated scintigraphy image are in the supplementary material, §S5.
@@ 3.1 Particle deposition in the lungs without airway constriction
Figure 2: Data from simulations in the unconstricted lung geometries. In (a)-(d), particle diameter is $4\mu\mathrm{m}$. (a) Deposited fraction within each conducting airway in one geometry, shown as a planar graph. Deposition at bifurcations is assigned to the parent airway. (b) Simulated scintigraphy plot for the same example. (c) Deposited fraction in each lobe, showing mean $\pm$ one standard deviation across the four geometries. Deposited fraction separated into conducting-airways and acinar deposition. (d) Deposited fraction in each generation, showing mean $\pm$ one standard deviation across the four geometries, with conducting-airways and acinar deposition indicated. Generational deposition from MPPD is also shown. (e) Total and conducting-airways deposition, versus particle diameter, with comparison to MPPD.
@@
To validate the model in healthy lung geometries, we compare deposition in the four lung geometries against results from MPPD (figure 2d,e). Deposition in each airway generation shows generally good agreement for $4\mu\mathrm{m}$ diameter particles (figure 2d), despite MPPD using a lung geometry based on different morphometric data. There is also good agreement with MPPD predictions of total deposition over a range of particle sizes, with only minor quantitative differences for particle diameters smaller than $0.1\mu\mathrm{m}$ or larger than $1\mu{\mathrm{m}}$. Higher conducting-airways deposition for large particles in our model is likely due to the central airways being narrower on average than in the MPPD lung geometry, causing higher impaction. We use different semi-empirical formulae than in MPPD to determine deposition rates, but despite this, agreement is still good.
The MPPD model has been validated against experimental data in people with healthy lungs [heyder1986deposition], showing good agreement in total deposition over a range of particle sizes from $0.01\mu\mathrm{m}$ to $10\mu\mathrm{m}$ diameter particles [asgharian2006prediction, asgharian_2022_mixing]. Good agreement (figure 2d,e) between our predictions of lung deposition and MPPD, therefore, also indicates that our model accurately captures lung deposition fractions across a wide range of particle sizes in healthy lungs. Comparison to the MPPD model provides validation in healthy lung geometries, but the MPPD model assumes a uniform ventilation so is not able to capture the effects of airway constriction or blockage on patterns of ventilation [asgharian2006prediction, hofmann2011modelling]. In the following sections, we demonstrate our model’s ability to predict heterogeneous patterns of ventilation and particle deposition after many airways in the lungs have been constricted.
@@ 3.2 Effects of varying airway constriction severity
Figure 3: Simulations in which all generation 12-15 airways in the left upper lobe are constricted with varying severities (fraction by which airway radii are reduced). In each simulation, the constriction severity is the same for all left upper lobe airways. (a) Change in deposition in the affected lobe (LU), and the other lobes, separated into central conducting-airways deposition, distal conducting-airways deposition and acinar deposition. (b) Lung Clearance Index (LCI), from MBW simulations in the same constricted geometries. (c) Change in individual airway deposition in the simulation with constriction severity 0.825.
@@
To explore how the severity of airway constriction impacts deposition, we first constrict all generation 12-15 airways in a single lobe, with a range of constriction severities. The effect of increasing constriction severity on deposition in the affected airways is non-monotonic (figure 3a). In these simulations, distal-airways deposition is not found to be significantly impacted by mild constrictions but, when the severity exceeds 0.4, it increases as impaction is significantly enhanced in the narrowed airways. As the severity is increased further beyond 0.675, the increased resistance of the affected airways causes flow to be reduced into the affected lobe to such a degree that distal deposition decreases rapidly. Increased impaction in narrowed airways has been highlighted previously in simpler models [martonen2003silico], but these have not captured the subsequent decrease as airway resistance is further increased and ventilation patterns are altered.
Central and acinar deposition in the affected lobe decrease as the constriction severity is increased and ventilation of the lobe is reduced (figure 3a). In the other lobes, deposition increases due to a higher proportion of the inhaled particles entering these lobes and the speed of air flow through them being faster. The strongest increase in deposition in the unconstricted lobes is in the central airways since impaction is the dominant deposition mechanism there, and this is enhanced by faster flow. By contrast, in the distal conducting airways, where sedimentation is the dominant mechanism, there is minimal change to deposition as shorter residence times for particles leads to reduced sedimentation, enough to largely offset any increased impaction. The changes in deposition that we predict in airways that are not themselves constricted cannot be reproduced by existing models such as MPPD [asgharian2001particle], which assume uniform ventilation. Figure 3(a) shows that when the left upper lobe airways are severely constricted, we predict up to an $80\%$ increase in central-airways deposition in the rest of the lungs, and almost no change in distal-airways deposition. This quantification of the relative impact on deposition in the central and distal airways suggests a mechanism for how localised airway constriction can increase the central-peripheral deposition ratio throughout the lungs, a measure that can be increased [de2010lung, virchow2018lung] in people with obstructive airway diseases, albeit with significant inter-subject variability in the experimental data.
As the constriction severity is increased in these simulations, other parameters such as particle size and breathing rate are kept constant, so as to isolate the impacts of airway constriction. The assumption of fixed tidal volume means that local changes in flow rate are solely due to redistribution of lobar ventilation. Some of the changes in regional deposition might be partially mitigated by adjusting these other parameters: for example, since impaction is weaker for smaller particles and at lower flow rates, the increase in central-airways deposition after applying severe constrictions may be smaller if the particle size or breathing rate was decreased. However, in the simulations shown in figure 3, we assume a typical flow rate for normal breathing and a particle size of $4\mu\mathrm{m}$, which is typical of the MMAD of many nebulised drugs.